OpenStudy (anonymous):

did i do something wrong here???? Evaluate the line integral ∫CF⋅dr, where F(x,y,z)=4sinxi+4cosyj−xzk and C is given by the vector function r(t)=t^6i−t^5j+t^4k , 0≤t≤1.

OpenStudy (anonymous):

did \[\int\limits_{0}^{1}<4\sin(t^6), 4\cos(-t^5), (t^6t^4)> \dot <6t^5, -5t^4, 4t^3>\] equals\[\int\limits_{0}^{1}(24t^5\sin(t^6)-20t^4\cos(-t^5)+4t^(13) dt \] plug that into a calculator and i get -1.24 and webwork says no. did i make a mistake somwhere?

OpenStudy (fibonaccichick666):

well, first, is it a dot product?

OpenStudy (anonymous):

yes. i didn't know how to do a dot in the thing, so i just wrote dot.

OpenStudy (anonymous):

it's r(t) dot r'(t)

ganeshie8 (ganeshie8):

try ` -1.813`

OpenStudy (anonymous):

yes, that works. what did i goof on?

ganeshie8 (ganeshie8):

look at ur z component after parameterization

OpenStudy (anonymous):

oops. -x, so -(t^6). i knew it was something simple like that. thanks!

OpenStudy (anonymous):

adhd makes my brain work faster than my pencil sometimes and i miss negatives....:D

OpenStudy (fibonaccichick666):

he found it faster than me

ganeshie8 (ganeshie8):

you're using wolfram right ? i think there is some wolfram applet for line integrals too http://www.wolframalpha.com/input/?i=%5Cint_0%5E1+%2824t%5E5%5Csin%28t%5E6%29-20t%5E4%5Ccos%28-t%5E5%29%2B4t%5E%2813%29%29++dt

OpenStudy (anonymous):

no, i was using integral-calculator.com. but i can use wolfram too

OpenStudy (fibonaccichick666):

Why not do it by hand?

OpenStudy (anonymous):

faster, fibonaccichick. i know how to integrate and some of these end up being really long integrals, so it's just faster once i get the integral set up.

OpenStudy (fibonaccichick666):

XD I know, but usually you are not allowed to use calculators on tests (at least for my calc 3) so it's good practice

OpenStudy (anonymous):

oh yeah, the tests have easier integrals than the webwork gives us though.

OpenStudy (anonymous):

for the test it'll be something like the integral of x +y or sin+cos instead of all this garbledy gook.

OpenStudy (anonymous):

lol

OpenStudy (fibonaccichick666):

ahhh ok, yea, If you have a TI 89 you can nicely do the integral on it :) but yes, this is nasty, very nasty. Side note, why in the j component did you keep a negative on the t^5? I thought vecotors were supposed to be wriiten I-j+k, so we would have to modify the sign? (It's been like 3 years since calc 3 for me)

OpenStudy (fibonaccichick666):

actually, no not that nasty: here it's just two u substitutions

ganeshie8 (ganeshie8):

yeah they are cookedup to evaluate nicely by hand

OpenStudy (anonymous):

yeah, we only have like 50 minutes to do 10 questions, so they make them shorter. :D thanks for all the help! onto question 8!

OpenStudy (fibonaccichick666):

it becomes \[\int _{0}^{1} 4sinu-4cosu-4t^{13}\]

OpenStudy (fibonaccichick666):

with implied du's and dt's

OpenStudy (anonymous):

it also leaves less room for mistakes for me. with my adhd, it's common for me to do stuff like that. as seen here. on a test once my teacher circled where i'd accidentally done 1-0=-9 and put a question mark next to it. he only took off 1 point cause even though i got the wrong answer cause of that goof, i still knew what i was doing. working to try and slow down my brain so i don't goof like that again

OpenStudy (fibonaccichick666):

well, on homework yes, but you will never get better without practice

OpenStudy (dan815):

hi